OpenStudy (anonymous):

compounded semiannually (n=2)?

OpenStudy (anonymous):

is that right?

OpenStudy (wolf1728):

I need a little more information.

OpenStudy (anonymous):

im asking if coumponded semiannually is equal to 2 algerbra 2 stuff

OpenStudy (wolf1728):

It all depends on the formula you are using. Compound rate = (1+ rate/(compounding periods)^compounding periods Compound rate = oh and what is the interest rater?

OpenStudy (anonymous):

How much will $600 amount to be if it is invested at 10% interest compounded semiannually (two times a year) for 4 years?

OpenStudy (anonymous):

im useing A(t)=P(1+ r/n)^nt formula

OpenStudy (anonymous):

is that right?

OpenStudy (wolf1728):

Is this only $600 invested once? Or is this an annuity in which the $600 is invested twice a year for 4 years?

OpenStudy (ranga):

Sorry I couldn't reply sooner. Was caught in long calculations with three others. In the compound interest formula, compounding semiannually means n = 2 quarterly means n = 4 yearly means n = 1 monthly means n = 12 daily means n = 365

OpenStudy (wolf1728):

The 10 % compounded semi-annually equals 10.25% Total = Principal * (1+rate)^4 Total = $600 * (1.1025)^4 Total = 600 * 1.4774554438 Total = 886.47

OpenStudy (ranga):

Yes. Reaper: The same result can be gotten using the formula given to you: A(t)=P(1+ r/n)^nt = 600(1 + 0.1/2)^(4*2) = $886.47

OpenStudy (wolf1728):

All right - great minds think alike!!

OpenStudy (ranga):

lol.

OpenStudy (anonymous):

haha thanks ranga better late then never right? lol